OpenStudy (anonymous):

Find V if r = 2 and L.A. = 12pi

OpenStudy (zzr0ck3r):

volume of what?

OpenStudy (zale101):

what's the shape?

OpenStudy (anonymous):

right circular cone

OpenStudy (zale101):

oh

OpenStudy (anonymous):

What is L.A.?

OpenStudy (anonymous):

lateral area

OpenStudy (zzr0ck3r):

so this is a cylinder?

OpenStudy (anonymous):

its a cone...

OpenStudy (zale101):

for the lateral area Height= sqrt(6^2 - 2^2) Height = sqrt(32) for the volume V = (1/3)*pi*r^2*h

OpenStudy (zale101):

also, the height can be written as 4 x sqrt(2) as an square root